### v2

# arXiv:2602.19591v2[cs.LG] 27 Feb 2026

Abstract
Small and Medium Enterprises (SMEs) constitute the vast majority of businesses globally and generate a significant portion of economic activity, yet systematically identifying high-potential SMEs remains an open challenge. We introduce SME-HGT, a Heterogeneous Graph Transformer framework that predicts which early-stage (Phase I) innovation grant awardees will advance to Phase II funding using exclusively public data. We construct a heterogeneous graph with 32,268 company nodes, 124 research topic nodes, and 13 funding agency nodes connected by approximately 99,000 edges across three semantic relation types. SME-HGT achieves an AUPRC of $0.621\pm 0.003$ on a temporally-split test set, outperforming an MLP baseline ($0.590\pm 0.002$) and R-GCN ($0.608\pm 0.013$) across five random seeds. At a screening depth of 100 companies, SME-HGT attains 89.6% precision with a $2.14\times$ lift over random selection. Our temporal evaluation protocol prevents information leakage, and our reliance on public data ensures reproducibility. These results demonstrate that relational structure among firms, research topics, and funding bodies provides meaningful signal for SME potential assessment, with implications for economic policymakers and early-stage investors globally.
I Introduction
Small and Medium Enterprises (SMEs) are the primary engine of innovation and economic growth globally. According to international economic development reports [18], SMEs account for the overwhelming majority of all businesses, employ a large segment of the private workforce, and generate a substantial portion of global economic activity. Identifying which of these firms have the highest growth potential is a longstanding challenge for economic analysts, investors, and grant program administrators alike.
National and regional innovation grant programs are major sources of early-stage technology funding for small businesses, distributing billions annually across various funding agencies [11]. These competitive programs typically operate in sequential phases: Phase I awards fund initial feasibility studies, while Phase II awards support full research and development. Progression from Phase I to Phase II is rigorously evaluated and serves as a meaningful proxy for a firm’s technical and commercial potential [12, 1]. Predicting which Phase I awardees will advance to Phase II is therefore both a practical policy question and a test case for data-driven firm assessment.
Current approaches to SME assessment rely predominantly on tabular features—financial metrics, patent counts, employee size—or manual expert review. These methods miss an important dimension: the relational structure in which firms are embedded. A company’s potential may be signaled not only by its own attributes but also by its connections to specific research domains, its funding relationships with sponsoring agencies, and its proximity to other successful firms in the innovation ecosystem.
Graph Neural Networks (GNNs) offer a natural framework for capturing such relational information. Recent work has demonstrated the power of GNNs for tasks ranging from social network analysis [8] to molecular property prediction [7]. Broader artificial intelligence methodologies are similarly transforming critical physical infrastructure, optimizing design and control strategies in power electronics and energy systems [4]. In the economic domain, heterogeneous graph methods have shown promise for firm-level prediction tasks [9], though prior work has typically relied on proprietary data sources.
In this paper, we make three contributions:
  1. 1.
    Public-data heterogeneous graph. We construct a heterogeneous graph of an innovation grant ecosystem using exclusively public database records, with three node types (companies, research topics, funding agencies) and three edge types capturing operational, funding, and co-topic relationships.
  2. 2.
    SME-HGT framework. We adapt the Heterogeneous Graph Transformer [9] for SME potential prediction, demonstrating that type-specific attention over heterogeneous relations outperforms both tabular baselines and homogeneous graph approaches.
  3. 3.
    Temporal evaluation protocol. We design a strict temporal split that prevents information leakage, using feature cutoffs, label horizons, and chronological train/validation/test partitions to simulate realistic deployment conditions.
II Related Work
II-A Graph Neural Networks for Economic Applications
Graph-based learning has been increasingly applied to economic and financial prediction tasks. Kipf and Welling [10] introduced Graph Convolutional Networks (GCNs) for semi-supervised node classification, and subsequent architectures including GAT [19] and GraphSAGE [8] have been applied to fraud detection, credit scoring, and supply chain analysis. Zhou et al. [21] provide a comprehensive survey of GNN architectures and applications.
In the context of firm-level prediction, GNNs can leverage corporate networks, ownership structures, and industry co-membership graphs to augment tabular features. However, most economic applications of GNNs operate on homogeneous graphs and rely on proprietary data. Our work addresses both limitations: we use a heterogeneous graph that explicitly models different entity types and their distinct relations, and we rely exclusively on publicly available grant program data.
II-B Innovation Grant Analytics
Innovation grant programs have been extensively studied in the economics literature. Lerner [11] conducted a foundational analysis showing that grant awardees exhibit significantly higher growth rates than matched non-awardees. Link and Scott [12] examined the role of these programs in bridging the “valley of death” between basic research and commercialization. Audretsch et al. [1] studied the relationship between early-stage funding and small firm innovation output. Toole and Czarnitzki [17] analyzed phase progression specifically in the biomedical sector.
Despite this rich literature, existing grant analyses are predominantly retrospective and descriptive. To our knowledge, no prior work has applied graph-based machine learning to predict Phase II progression using the relational structure of the innovation ecosystem.
II-C Heterogeneous Graph Neural Networks
Real-world graphs frequently contain multiple node and edge types. Several architectures have been proposed for heterogeneous graph learning. Schlichtkrull et al. [15] introduced R-GCN with relation-specific weight matrices. Wang et al. [20] proposed Heterogeneous Graph Attention Network (HAN) using metapath-based aggregation, building on the metapath concept from Sun et al. [16] and Dong et al. [5]. Most relevant to our work, Hu et al. [9] developed the Heterogeneous Graph Transformer (HGT), which uses type-specific linear projections and multi-head attention without requiring predefined metapaths.
We adopt HGT as our primary architecture due to its flexibility in handling arbitrary heterogeneous schemas and its ability to learn type-dependent attention patterns, which we hypothesize are important for capturing the distinct roles of companies, research topics, and funding agencies.
III Methodology
III-A Data
We use publicly available data from an open innovation awards database, which contains records of public grants made by participating funding bodies. The raw dataset comprises 207,726 award records spanning multiple decades. Each record includes the company name, award amount, award year, program phase (I or II), awarding agency, and research topic code.
We apply the following data cleaning pipeline: (1) standardize company names by converting to uppercase, removing legal suffixes, and collapsing whitespace for entity resolution; (2) parse phase numbers from heterogeneous string formats; (3) extract research topic prefixes and agency identifiers; and (4) remove records with missing critical fields. After entity resolution, the cleaned dataset contains 32,268 unique companies.
III-B Graph Construction
We construct a heterogeneous graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with three node types and three semantic edge types. Reverse edges are automatically added for bidirectional message passing, yielding six directed edge types and approximately 99,000 total edges. Table I summarizes the graph schema.
TABLE I: Heterogeneous Graph Schema
Node Type Count Features
Company 32,268 7
Topic 124 2
FundingAgency 13 3
Edge Types
(Company, operates_in, Topic)
(Company, awarded_by, FundingAgency)
(Company, co_topic, Company)
Company features (7 dimensions) characterize each firm’s Phase I history: log-transformed total Phase I funding, log Phase I award count, number of unique funding agencies (agency diversity), years active in the program, log average Phase I award size, topic diversity (number of unique research topics), and a normalized award recency feature. All features are min-max normalized to $[0,1]$.
Topic features (2 dimensions) characterize each research area: log number of participating companies and log total awards within the topic, both normalized.
Funding agency features (3 dimensions): log award count, log total funding amount, and log average award size, all normalized.
Edge construction. Each company is linked to its primary (modal) research topic via an operates_in edge and to its primary funding body via an awarded_by edge. Companies sharing the same primary research topic are connected by co_topic edges, with a cap of 50 edges per topic group and 20 edges per node to manage density in large topic clusters.
Temporal integrity. All company features are computed using only Phase I awards prior to the feature cutoff date (January 1, 2018). Phase II information is never included in node features, ensuring strict separation between inputs and prediction targets.
III-C Label Design
We define the prediction target as a binary variable: whether a Phase I awardee receives a Phase II award within a 5-year horizon of their first Phase I award. Formally, for company $c$:
$y_{c}=\mathbb{1}\!\left[\exists\text{ Phase II award for }c\text{ with year}\leq t^{(1)}_{c}+5\right]$ (1)
where $t^{(1)}_{c}$ denotes the year of company $c$’s first Phase I award.
We partition companies into temporally non-overlapping sets based on their first Phase I award year:
  • Train: first Phase I year <2018<2018
  • Validation: $2018\leq$ first Phase I year <2020<2020  (n=2,353n=2{,}353)
  • Test: $2020\leq$ first Phase I year <2022<2022  (n=2,689n=2{,}689)
Companies with first Phase I year $\geq 2022$ are excluded due to insufficient observation window for the 5-year label horizon. The test set positive rate is 41.8% ($n_{+}=1{,}124$), reflecting moderate class balance.
III-D SME-HGT Architecture
Our model, SME-HGT, consists of three components: type-specific input projection, heterogeneous graph transformer layers, and a classification head. Figure 1 illustrates the overall architecture.
[…]
Figure 1: SME-HGT architecture. Input features are projected per node type to a shared dimension, then refined through three HGT layers with residual connections. Only company node embeddings are passed to the classifier.
Input projection. Each node type $\tau\in\{\text{Company},\text{Topic},\text{FundingAgency}\}$ receives a dedicated linear transformation $\mathbf{h}^{(0)}_{v}=\mathbf{W}_{\tau}\mathbf{x}_{v}+\mathbf{b}_{\tau}$, mapping heterogeneous input features to a shared hidden dimension $d=128$.
HGT layers. We stack $L=3$ Heterogeneous Graph Transformer [9] layers. Each layer computes type-dependent attention weights across all edge types using $H=4$ attention heads. For a target node $t$ receiving messages from source node $s$ via relation $r$:
$\alpha_{s,r,t}=\text{Softmax}_{s\in\mathcal{N}_{r}(t)}\!\left(\frac{(\mathbf{W}^{Q}_{\tau(t)}\mathbf{h}_{t})^{\top}(\mathbf{W}^{K}_{\tau(s),r}\mathbf{h}_{s})}{\sqrt{d/H}}\right)$ (2)
[…]
Each HGT layer is followed by a residual connection, per-type layer normalization [2], and dropout ($p=0.2$).
Classification head. After $L$ layers of message passing, company node representations are passed through a two-layer MLP classifier:
$\hat{y}_{c}=\text{Linear}_{64\to 2}\!\left(\text{Dropout}\!\left(\text{ReLU}\!\left(\text{Linear}_{128\to 64}(\mathbf{h}^{(L)}_{c})\right)\right)\right)$ (4)
producing logits for binary classification. Topic and agency node embeddings are not directly classified; they serve as context through message passing.
III-E Training Procedure
We optimize all models using AdamW [14] with learning rate $10^{-3}$ and weight decay $5\times 10^{-4}$. The learning rate follows a linear warmup over 10 epochs followed by cosine annealing [13] over the remaining epochs. Training proceeds for a maximum of 200 epochs with early stopping based on validation loss (patience =30=30). Gradient norms are clipped at 1.0. The loss function is binary cross-entropy, appropriate given the near-balanced class distribution in our dataset. All experiments are implemented using PyTorch Geometric [6].
IV Experiments
IV-A Experimental Setup
Baselines. We compare SME-HGT against two baselines:
  • MLP: A 3-layer feedforward network operating on company node features only, ignoring all graph structure. Each layer applies Linear $\to$ ReLU $\to$ BatchNorm $\to$ Dropout with hidden dimension 128. This baseline isolates the contribution of graph structure by using identical features.
  • R-GCN [15]: A relational graph convolutional network using SAGEConv [8] layers converted to heterogeneous form via per-type input projections. It uses 2 message-passing layers with residual connections and LayerNorm, followed by the same classification head.
All models share the same hidden dimension (128), dropout rate (0.2), and training hyperparameters. Model sizes are: SME-HGT 729,863 parameters; R-GCN 340,674; MLP 18,178. All results are averaged over 5 random seeds $\{42,123,456,789,1024\}$.
Metrics. We report Area Under the Precision-Recall Curve (AUPRC) as the primary metric, following recommendations for evaluation under moderate class imbalance [3]. We additionally report AUROC, F1 score at the optimal threshold, Precision@$K$ for $K\in\{100,500,1000\}$, and Lift@$K$ defined as Precision@$K$ divided by the base positive rate.
IV-B Main Results
Table II presents the primary experimental results on the held-out test set.
TABLE II: Test Set Results (mean $\pm$ std over 5 seeds)
Model AUPRC AUROC F1
[…]
SME-HGT achieves the highest AUPRC of $0.621\pm 0.003$, a 3.1 percentage point (pp) improvement over the MLP baseline and 1.3 pp over R-GCN. The AUROC gap is larger: +4.3+4.3 pp over MLP and +0.5+0.5 pp over R-GCN, indicating improved ranking quality across all operating thresholds. F1 scores at optimal thresholds are comparable ($\sim$0.591), reflecting similar behavior at each model’s best binary decision boundary.
IV-C Ranking Performance
Table III presents Precision@$K$ and Lift@$K$ metrics, which are most relevant for practical screening applications where analysts review a fixed number of top-ranked candidates.
TABLE III: Ranking Metrics on Test Set (mean over 5 seeds)
Precision@$K$ Lift@$K$
Model @100 @500 @1K @100 @500 @1K
MLP .896 .675 .531 2.14 1.61 1.27
R-GCN .886 .698 .559 2.12 1.67 1.34
HGT .896 .712 .569 2.14 1.70 1.36
At the top-100 screening depth, all three models achieve similar precision ($\sim$89–90%), indicating that the highest-confidence predictions are robust across architectures. However, as the screening depth increases, the advantage of graph-based methods becomes evident. SME-HGT achieves 71.2% Precision@500 compared to 67.5% for MLP (+3.7+3.7 pp) and 69.8% for R-GCN (+1.4+1.4 pp). This pattern intensifies at $K=1000$, where SME-HGT leads MLP by 3.8 pp. All models substantially outperform random selection (Lift >1.0>1.0), with SME-HGT maintaining the highest lift at every screening depth.
IV-D Analysis
Graph structure improves discrimination. The consistent AUPRC gap between graph-based methods (HGT, R-GCN) and the non-graph MLP baseline confirms that relational structure in the innovation ecosystem carries predictive signal beyond what tabular company features alone capture. The company-topic and company-agency relations, as well as co-topic links between firms, each contribute information about a firm’s context.
HGT offers the best stability–performance tradeoff. While R-GCN achieves competitive mean performance, its AUPRC standard deviation ($\pm 0.013$) is over four times that of SME-HGT ($\pm 0.003$). This suggests R-GCN is more sensitive to random initialization, likely because its relation-specific weight matrices interact with the relatively small number of edge types (3) in our graph. The MLP exhibits the lowest variance ($\pm 0.002$) but at the cost of substantially lower mean performance.
Convergence behavior. SME-HGT converges in 23–41 epochs across seeds (mean training time $\sim$149 s), while R-GCN requires 26–77 epochs ($\sim$31 s) and MLP 50–56 epochs ($\sim$8 s). Despite the higher per-epoch cost of attention-based message passing, all training times remain within practical bounds for an offline screening application.
Moderate overall discrimination. We note that absolute AUROC values (0.60–0.65) indicate moderate rather than strong discriminative power. This reflects the inherent difficulty of predicting Phase II outcomes—which depend on proposal quality, reviewer composition, and evolving funding priorities—from structural and historical features alone. Nevertheless, the ranking metrics demonstrate substantial practical value: screening the top 100 candidates yields nearly $90\%$ precision, more than double the base rate.
V Discussion
V-A Policy Implications
Our results have direct relevance for global grant program administration and small business policy. A screening tool achieving 89.6% precision among its top 100 predictions could substantially reduce the manual review burden for program managers. At a base rate of 41.8%, random selection would yield approximately 42 successful firms out of 100 reviewed; SME-HGT identifies approximately 90. This $2.14\times$ lift translates to more efficient allocation of expert review resources and potentially earlier identification of high-potential firms for targeted technical assistance.
The public-data-only design is a deliberate choice: any jurisdiction with structured grant or subsidy data can construct similar heterogeneous graphs to assess participant potential without requiring proprietary databases. This makes our approach replicable for analogous programs across various national innovation systems.
V-B Limitations
Several limitations warrant discussion. First, our graph relies exclusively on historical grant program data; incorporating additional public sources such as patent filings or local economic data could strengthen both node features and graph connectivity. Second, we observe only Phase I survivors—companies that already won competitive grants—introducing selection bias relative to the broader SME population. Third, the moderate overall discrimination ($\text{AUROC}\approx 0.65$) suggests that important predictive signals—proposal text quality, team composition, technology readiness level—are not captured in our current structural features. Fourth, our entity resolution relies on rule-based string matching, which may fail on name variants or subsidiaries not covered by our normalization rules. Finally, our evaluation uses a single temporal split; additional evaluation across different time windows would strengthen robustness claims.
V-C Future Work
Several directions merit exploration: (1) incorporating textual features from award abstracts via language model embeddings to capture proposal quality signals; (2) adding temporal dynamics through time-aware graph architectures such as evolving GCNs; (3) conducting ablation studies to quantify the marginal contribution of each node and edge type; (4) extending to downstream prediction targets such as patent filing, revenue growth, or acquisition events; and (5) developing an interactive dashboard for program managers to explore model predictions alongside company profiles.
VI Conclusion
We presented SME-HGT, a Heterogeneous Graph Transformer framework for identifying high-potential small businesses using public grant program data. By constructing a heterogeneous graph capturing relationships among 32,268 companies, 124 research topics, and 13 funding agencies, we demonstrated that relational structure provides meaningful predictive signal beyond tabular features alone. SME-HGT achieves an AUPRC of 0.621 on temporally-split test data, outperforming both MLP (+3.1+3.1 pp) and R-GCN (+1.3+1.3 pp) baselines while maintaining low variance across seeds. At a practical screening depth of 100 candidates, the model attains 89.6% precision—a $2.14\times$ improvement over random selection. Our strict temporal evaluation protocol and exclusive reliance on public data ensure that these results reflect realistic deployment conditions. This work illustrates the potential of heterogeneous graph neural networks for policy-relevant prediction tasks in the global SME innovation ecosystem.
Acknowledgments
Data sourced from public open-data grant portals. Computational experiments were conducted using PyTorch and PyTorch Geometric.